OpenStudy (anonymous):

Which of the following is the slope between the points (2, 3) and (-3, 0)? -5/3 5/3 -3/5 3/5

OpenStudy (diyadiya):

Slope of a line passing through two points (x1,y1) & (x2,y2) is given by the formula \[\large Slope = \frac{y2-y1}{x2-x1}\]

OpenStudy (diyadiya):

Can you try to plug in the values of x1,y1,x2 & y2 and find the slope?

OpenStudy (anonymous):

option D is right answer

OpenStudy (diyadiya):

(x1,y1) = (2,3) (x2,y2) = (-3,0) x1=2 y1=3 x2=-3 y2=0 \[\large Slope = \frac{y2-y1}{x2-x1}= \frac{0 - 3}{-3-2}\]